jim can fill a pool carrying buckets of water in 30 mins. sue can do the same job in 45 mins. tony can do the same job in an hour and a half. how quickly can all 3 fill the pool together?
Answer by richwmiller(17219)

x/30+x/45+x/90=1
3x+2x+x=90
6x=90
x=15 minutes for all three
3/6+2/6+1/6=1